OpenStudy (anonymous):

just a quick question. When trying to determine the zeros of a function and you are finding the negative real zeros, how exactly to you write the function f(-x). Like what i mean is how do you know which signs need to be changed?

OpenStudy (jdoe0001):

ahemmm, well, you'd want to change all "x" to "-x" and it makes it easier to avoid confusion if you enclose them in parentheses

OpenStudy (anonymous):

For example with these three equations i have. \[f(x)=x^3-3x^2-16x+48\] \[f(x)=x^3-12x^2+28x-9\] \[f(x)=x^4-x^3+7x^2-9x-18\]

OpenStudy (jdoe0001):

say f(x) = 2x^3+x^2 f( -x ) = 2(-x)^3 + (x)^2 -x * -x * -x = (-x)^3 -x * -x * -x = -x^3 -x * -x = (-x)^2 -x * -x = x^2 f( -x ) = -2x^3 + x^2

OpenStudy (jdoe0001):

the rule goes, if the exponent is ODD, on the -x it will change sign due to the unevenness of the sign multiplication if the exponent is EVEN, it will end up as positive same reason, due to the evenness of the sign multiplication - * - * - = - -* - = +

OpenStudy (anonymous):

oh okay thank you(:

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

One thing. for like the first equation i put, \[f(x)=x^3-3x^2-16x+48\] does it become \[f(-x)==x^3+3x^2+16x+48\]

OpenStudy (anonymous):

And the second one from \[f(x)=x^3-12x^2+28x-9\] to \[f(-x)=-x^3+12x^2-28x-9\]

OpenStudy (anonymous):

And the last one from \[f(x)=x^4-x^3+7x^2-9x-18\] to \[f(-x)=x^4+x^3+7x^2+9x-18\]

OpenStudy (anonymous):

@jdoe0001 ?